# From smooth dynamical twists to twistors of quantum groupoids

### v1

@@ Abstract.
Consider a Lie subalgebra ${\mathfrak{l}}\subset{\mathfrak{g}}$ and an ${\mathfrak{l}}$-invariant open submanifold $V\subset{\mathfrak{l}}^{\ast}$. We demonstrate that any smooth dynamical twist on $V$, valued in $U({\mathfrak{g}})\otimes U({\mathfrak{g}})\llbracket\hbar\rrbracket$, establishes a twistor on the associated quantum groupoid when combined with the Gutt star product on the cotangent bundle $T^{\ast}L$ of a Lie group $L$ that integrates ${\mathfrak{l}}$. This result provides a framework for constructing equivariant star products from dynamical twists on those Poisson homogeneous spaces arising from nondegenerate polarized Lie algebras, leveraging the structure of twistors of quantum groupoids.
@@ 1. Introduction
One of Drinfeld’s significant result asserts that solutions to the quantum Yang-Baxter equation give rise to quantum groups [Drinfeld-QuantumGroup, Drinfeld-HopfYangBaxter]. Similarly, solutions to the quantum dynamical Yang-Baxter equation (QDYBE), also known as the Gervais-Neveu-Felder equation [Gervais-Neveu, Felder], lead to the formation of dynamical quantum groups [Etingof-Varchenko-Exchange, Etingof-Varchenko]. We recall a useful method for finding solutions to the QDYBE established in [Babelon-Bernard-Billey]. Let ${\mathfrak{g}}$ be a finite-dimensional Lie algebra, and $\eta$ an abelian subalgebra of ${\mathfrak{g}}$. Consider a smooth function $F$ defined on an open submanifold $V$ of $\eta^{\ast}$, taking values in $U({\mathfrak{g}})\otimes U({\mathfrak{g}})\llbracket\hbar\rrbracket$, where $F\in C^{\infty}(V)\llbracket\hbar\rrbracket\otimes U({\mathfrak{g}})\otimes U({\mathfrak{g}})$ and $\hbar$ is a formal variable. Assume further that $F$ is of the form $F=1\otimes 1\otimes 1+O(\hbar)$ and is $\eta$-invariant. The function $\mathsf{R}:=F_{21}^{-1}\cdot F$ is a solution to the associated QDYBE over the abelian base $\eta^{\ast}$ if and only if $F$ satisfies a shifted cocycle condition. This function $F$ is referred to as a smooth dynamical twist [Babelon-Bernard-Billey, Etingof-Nikshych-root, Etingof-Varchenko-Exchange].
@@
The formation of an elliptic quantum group, as introduced by Felder [Felder-Elliptic, Felder], originates from a smooth dynamical twist $F$. This quantum group is characterized as a family of quasi-Hopf algebras, denoted by $H_{\lambda}:=(U({\mathfrak{g}})\llbracket\hbar\rrbracket,\Delta_{\lambda})$, which are parameterized by $\lambda\in\eta^{\ast}$. The co-product $\Delta_{\lambda}$ is defined as $F^{-1}(\lambda)\Delta F(\lambda)$, resulting in a family of co-products on $U({\mathfrak{g}})\llbracket\hbar\rrbracket$ that are not necessarily coassociative. An important advancement in this area is presented in [Xu-R-matrices], where the algebra ${\mathcal{H}}={\mathcal{D}}\otimes U({\mathfrak{g}})\llbracket\hbar\rrbracket$ is enriched by ${\mathcal{D}}$, the algebra of smooth differential operators on $V$. This enrichment allows ${\mathcal{H}}$ to possess a quantum groupoid structure, as defined in [Xu-QuantumCNRS]. Quantum groupoids, a special class of Hopf algebroids, serve as a unification of quantum groups and star products [Xu-Quantum]. In this context, the role of $F$ is supplanted by ${\mathcal{F}}:=F\cdot\Theta_{0}$, where $\Theta_{0}$ is expressed as $\sum_{i}\hbar\frac{\partial}{\partial\lambda_{i}}\otimes l_{i}$. This ${\mathcal{F}}$, referred to as a twistor [Xu-R-matrices] of ${\mathcal{H}}$, transforms ${\mathcal{H}}$ into a new quantum groupoid, ${\mathcal{H}}_{{\mathcal{F}}}$, known as the dynamical quantum groupoid. Notably, the co-product in ${\mathcal{H}}_{{\mathcal{F}}}$ is co-associative and extends the co-product $\Delta_{\lambda}$ of $H_{\lambda}$ in a natural manner. Another construction of Hopf algebra alike objects associated with solutions of the QDYBE, known as the dynamical quantum groups, was developed in [Etingof-Varchenko, Etingof-Varchenko-Exchange, Etingof-Nikshych-root].
@@
Classical limits of smooth dynamical twists are triangular classical dynamical $r$-matrices which satisfy the classical dynamical Yang-Baxter equation (CDYBE) [Xu-Nonabelian]. A more general shifted-cocycle equation for dynamical twists with an associator term is proposed in [Enriquez-Etingof1], see [Enriquez-Etingof-Marshall] on further developments. The quantization of the Alekseev-Meinrenken dynamical $r$-matrices [Alekseev-Meinrenken], which are non-triangular and are over the nonabelian base ${\mathfrak{g}}^{\ast}$, are constructed in such setting [Enriquez-Etingof1]. The classical limits of dynamical twists in [Enriquez-Etingof1] are solutions of the modified CDYBE [LX]. In this note, we only consider the trivial associator case. For discussions of Hopf algebroids and quantum groupoids in other fields like knot and $3$-manifold invariants, dg geometry, and shifted Poisson geometry, we refer the reader to [Chen-Xiang-Xu, Kalmykov-Safronov, Nikshych-Turaev-Vainerman].
Consider a Lie group $L$ that integrates the Lie algebra $\mathfrak{l}$. Gutt [Gutt] developed a specific star product, now referred to as the Gutt star product, on the cotangent bundle $T^{\ast}L\cong\mathfrak{l}^{\ast}\times L$ of $L$. This star product, when regarded as a bi-differential operators on $V\times L\subset{\mathfrak{l}}^{\ast}\times L$, can be represented by an element
@@
[…]
@@
  • The left bottom-to-top-arrow is due to [Xu-Nonabelian]*Corollary 3.6. In fact, this arrow appeared even earlier than the introduction of quantum groupoids in [Xu-QuantumCNRS] in the setting $\eta\subset{\mathfrak{g}}$ being abelian.
  • The left top-to-bottom-arrow is pointed out in the proof of  [Alekseev-Calaque]*Proposition 1.3. Thus, the left vertical correspondence is one-to-one.
  • The right vertical correspondence is a bijection proved in [Xu-QuantumCNRS]. It is further shown op. cit. that star products can be equivalently described in terms of twistors of quantum groupoids.
@@
This note is structured to guide the reader through a comprehensive exploration of the topic. In Section 2, we revisit fundamental concepts related to Hopf algebroids, twistors, quantum groupoids, and dynamical twists. Section 3 is dedicated to state and prove Theorem A. Finally, in Section 4, we apply Theorem A to the Enriquez-Etingof formal dynamical twists, which arise from nondegenerate polarized Lie algebras [Enriquez-Etingof2]. This application offers a novel perspective on equivariant star products on the associated Poisson homogeneous spaces, approached from the standpoint of quantum groupoids.
@@
  • The notation ${\mathbb{K}}\llbracket\hbar\rrbracket$ stands for the algebra of formal power series over ${\mathbb{K}}$.
@@ 2.1. Hopf algebroids and twistors
    $\Delta\colon{\mathcal{H}}\rightarrow{\mathcal{H}}\otimes_{R}{\mathcal{H}}$, called the coproduct, is an $(R\,,R)$-bimodule morphism, satisfying $\Delta(1)=1\otimes 1$ and
@@
    $\epsilon\colon{\mathcal{H}}\rightarrow R$, called the counit, is an $(R\,,R)$-bimodule map, satisfying $\epsilon(1)=1$ and
    $(\epsilon\otimes_{R}{\mathrm{id}})\circ\Delta=({\mathrm{id}}\otimes_{R}\epsilon)\circ\Delta={\mathrm{id}}\,,$
    Here we have used the identification $R\otimes_{R}{\mathcal{H}}\simeq{\mathcal{H}}\otimes_{R}R\simeq{\mathcal{H}}$.
@@
  • We follow [Lu] to add Condition (6) in our definition, which is however not required in [Xu-Quantum]. In fact, we need this condition to endow $R$ with a left ${\mathcal{H}}$-module structure :×$\rhd\colon{\mathcal{H}}\times R\rightarrow R$ defined by
@@
    $R_{{\mathcal{F}}}:=(R\,,\ast_{{\mathcal{F}}})$ is an associative ${\mathbb{K}}$-algebra with unit $1$.
@@ 2.2. Quantum groupoids and star products
Given a smooth manifold $M$, let $R$ be a completed associative ${\mathbb{K}}\llbracket\hbar\rrbracket$-algebra with unit such that $R/\hbar R=C^{\infty}(M)$.
@@
A quantum universal enveloping algebroid (or QUE algebroid), also called a quantum groupoid, is a topological Hopf algebroid $({\mathcal{H}}\,,\alpha\,,\beta\,,\cdot_{\hbar}\,,\Delta\,,\epsilon)$ over $R$ (i.e., ${\mathcal{H}}$ is a completed ${\mathbb{K}}\llbracket\hbar\rrbracket$-module, all maps $\alpha$, $\beta$, $\cdot$, $\Delta$, $\epsilon$ are continuous with respect to the $\hbar$-adic topology), such that the induced Hopf algebroid ${\mathcal{H}}/\hbar{\mathcal{H}}$ over $C^{\infty}(M)$ is isomorphic to the standard Hopf algebroid $U(A)$ of a Lie algebroid $A$ over $M$.
@@
The Hopf algebroid structure on $U(TM)$ over $C^{\infty}(M)$ extends naturally to a quantum groupoid structure on ${\mathcal{H}}:=U(TM)\llbracket\hbar\rrbracket$ over $R:=C^{\infty}(M)\llbracket\hbar\rrbracket$ Consider a formal power series in $\hbar$
@@
for all $f,g\in C^{\infty}(M)\llbracket\hbar\rrbracket$ is a star product on $M$.
@@
Let $({\mathfrak{l}},[-,-]_{{\mathfrak{l}}})$ be an $N$-dimensional ${\mathbb{K}}$-Lie algebra. Denote by ${\mathfrak{l}}_{\hbar}:={\mathfrak{l}}\llbracket\hbar\rrbracket={\mathfrak{l}}\otimes{\mathbb{K}}\llbracket\hbar\rrbracket$ the Lie algebra whose Lie bracket is ${\mathbb{K}}\llbracket\hbar\rrbracket$-linear and satisfies $[X\,,Y]_{{\mathfrak{l}}_{\hbar}}=\hbar[X\,,Y]_{{\mathfrak{l}}}$, for all $X\,,Y\in{\mathfrak{l}}$. Consider the universal enveloping algebra $U({\mathfrak{l}}_{\hbar})$ of ${\mathfrak{l}}_{\hbar}$ over ${\mathbb{K}}\llbracket\hbar\rrbracket$. Note that
$U({\mathfrak{l}}_{h})/\hbar U({\mathfrak{l}}_{h})\simeq S({\mathfrak{l}})\cong\mathrm{Pol}({\mathfrak{l}}^{\ast}),$
@@
for all $f,g\in C^{\infty}({\mathfrak{l}}^{\ast})$, where {,}$\{-,-\}_{\pi_{{\mathfrak{l}}^{\ast}}}$ is the Poisson bracket induced by the linear Poisson structure $\pi_{{\mathfrak{l}}^{\ast}}$ on ${\mathfrak{l}}^{\ast}$, and $B_{k}$’s are a family of bi-differential operators on ${\mathfrak{l}}^{\ast}$ (see [Dito, Kathotia, Gutt]).
@@
Let ${\mathfrak{g}}$ be a Lie algebra such that ${\mathfrak{l}}\subset{\mathfrak{g}}$ is a Lie subalgebra. Let $G$ and $L$ be two Lie groups integrating ${\mathfrak{g}}$ and ${\mathfrak{l}}$, respectively. For any $l\in{\mathfrak{l}}\subset{\mathfrak{g}}$, denote by $\overset{\rightarrow}{l}$ the left invariant vector field on $L$ or on $G$ generated by $l$.
@@
    for some $F\in C^{\infty}(V)\llbracket\hbar\rrbracket\otimes U({\mathfrak{g}})\otimes U({\mathfrak{g}})$.
@@
An $F$ subject to the equations (2.11) and (2.12) is called a dynamical twist if it is also ${\mathfrak{l}}$-invariant [Enriquez-Etingof2, Enriquez-Etingof1]. Thus, a right $L$-invariant compatible star product on $V\times G$ determines a dynamical twist. For the converse, see Theorem 3.1 and the explanation on its relations with some existing works in the next section.

### v2

@@ Abstract.
Consider a Lie subalgebra ${\mathfrak{l}}\subset{\mathfrak{g}}$ and an $\mathfrak{l}$}-invariant open submanifold $V\subset{\mathfrak{l}}^{\ast}$. We demonstrate that any smooth dynamical twist on $V$, valued in $U({\mathfrak{g}})\otimes U({\mathfrak{g}})\llbracket\hbar\rrbracket$, establishes a twistor on the associated quantum groupoid when combined with the Gutt star product on the cotangent bundle $T^{\ast}L$ of a Lie group $L$ that integrates $\mathfrak{l}$}. This result provides a framework for constructing equivariant star products from smooth dynamical twists on those Poisson homogeneous spaces arising from nondegenerate polarized Lie algebras, leveraging the structure of twistors of quantum groupoids.
@@ 1. Introduction
One of Drinfeld’s significant result asserts that solutions to the quantum Yang-Baxter equation give rise to quantum groups [Drinfeld-QuantumGroup, Drinfeld-HopfYangBaxter]. Similarly, solutions to the quantum dynamical Yang-Baxter equation (QDYBE), also known as the Gervais-Neveu-Felder equation [Gervais-Neveu, Felder], lead to the formation of dynamical quantum groups [Etingof-Varchenko-Exchange, Etingof-Varchenko]. We recall a useful method for finding solutions to the QDYBE established in [Babelon-Bernard-Billey]. Let $\mathfrak{g}$} be a finite-dimensional Lie algebra, and $\eta$ an abelian subalgebra of $\mathfrak{g}$}. Consider a smooth function $F\in C^{\infty}(V)\llbracket\hbar\rrbracket\otimes U({\mathfrak{g}})\otimes U({\mathfrak{g}})$ defined on an open submanifold $V$ of $\eta^{\ast}$, and valued in $U({\mathfrak{g}})\otimes U({\mathfrak{g}})\llbracket\hbar\rrbracket$, where $\hbar$ is a formal variable. Assume further that $F$ is of the form $F=1\otimes 1\otimes 1+O(\hbar)$ and is $\eta$-invariant. The function $\mathsf{R}:=F_{21}^{-1}\cdot F$ is a solution to the associated QDYBE over the abelian base $\eta^{\ast}$ if and only if $F$ satisfies a shifted cocycle condition. This function $F$ is referred to as a smooth dynamical twist [Babelon-Bernard-Billey, Etingof-Nikshych-root, Etingof-Varchenko-Exchange].
@@
The formation of an elliptic quantum group, as introduced by Felder [Felder-Elliptic, Felder], originates from a smooth dynamical twist $F$. This quantum group is characterized as a family of quasi-Hopf algebras, denoted by $H_{\lambda}:=(U({\mathfrak{g}})\llbracket\hbar\rrbracket,\Delta_{\lambda})$, which is parameterized by $\lambda\in\eta^{\ast}$. The coproduct $\Delta_{\lambda}$ is defined as $F^{-1}(\lambda)\Delta F(\lambda)$, resulting in a family of coproducts on $U({\mathfrak{g}})\llbracket\hbar\rrbracket$ that are not necessarily coassociative. An important advancement in this area is presented in [Xu-R-matrices], where the algebra ${\mathcal{H}}={\mathcal{D}}\otimes U({\mathfrak{g}})\llbracket\hbar\rrbracket$ is enriched by the algebra $\mathcal{D}$}, the algebra of smooth differential operators on $V$. This enrichment allows $\mathcal{H}$} to possess a quantum groupoid structure, as defined in [Xu-QuantumCNRS]. Quantum groupoids, a special class of Hopf algebroids, serve as a unification of quantum groups and star products [Xu-Quantum]. In this context, the role of $F$ is supplanted by ${\mathcal{F}}:=F\cdot\Theta_{0}$, where $\Theta_{0}$ is expressed as $\exp(\sum_{i}\hbar\frac{\partial}{\partial\lambda_{i}}\otimes l_{i})$. This $\mathcal{F}$}, referred to as a twistor [Xu-R-matrices] of $\mathcal{H}$}, transforms $\mathcal{H}$} into a new quantum groupoid $\mathcal{H}$$_{{\mathcal{F}}}$, known as the dynamical quantum groupoid. Notably, the coproduct in $\mathcal{H}$$_{{\mathcal{F}}}$ is coassociative and extends the coproduct $\Delta_{\lambda}$ of $H_{\lambda}$ in a natural manner. Another construction of Hopf algebra alike objects associated with solutions of the QDYBE, known as the dynamical quantum groups, was developed in [Etingof-Varchenko, Etingof-Varchenko-Exchange, Etingof-Nikshych-root, Felder-Elliptic].
@@
Classical limits of smooth dynamical twists are triangular classical dynamical $r$-matrices which satisfy the classical dynamical Yang-Baxter equation (CDYBE) [Xu-Nonabelian]. A more general shifted-cocycle equation for dynamical twists with an associator term is proposed in [Enriquez-Etingof1], we refer to [Enriquez-Etingof-Marshall] on further developments. The quantization of the Alekseev-Meinrenken dynamical $r$-matrices [Alekseev-Meinrenken], which are non-triangular and are over the nonabelian base $\mathfrak{g}$$^{\ast}$, is constructed in such setting [Enriquez-Etingof1]. The classical limits of dynamical twists in [Enriquez-Etingof1] are solutions of the modified CDYBE [LX]. In this note, we only consider the trivial associator case. For discussions of Hopf algebroids and quantum groupoids in other fields like invariants of knots and $3$-manifold invariants, dg geometry, and shifted Poisson geometry, we refer the reader to [Chen-Xiang-Xu, Kalmykov-Safronov, Nikshych-Turaev-Vainerman].
Consider a Lie group $L$ that integrates the Lie algebra $\mathfrak{l}$. Gutt [Gutt] developed a specific star product, now referred to as the Gutt star product, on the cotangent bundle $T^{\ast}L\cong\mathfrak{l}^{\ast}\times L$ of $L$. A normal ordering version of this star product [Xu-Nonabelian], when regarded as a bi-differential operator on $V\times L\subset{\mathfrak{l}}^{\ast}\times L$, can be represented by an element
@@
$\Bigg\{\!\!\!\!\!\begin{aligned} &&\text{smooth dynamical twists}\qquad\\ &&F\in C^{\infty}(V)\llbracket\hbar\rrbracket\otimes U({\mathfrak{g}})\otimes U({\mathfrak{g}})\end{aligned}\Bigg\}$$\Bigg\{\text{twistors ${\mathcal{F}}$ of ${\mathcal{H}}={\mathcal{D}}\otimes U({\mathfrak{g}})\llbracket\hbar\rrbracket$}\Bigg\}$$\Bigg\{\!\!\!\!\begin{aligned} &&\text{right $L$-invariant compatible}\quad\quad\\ &&\text{star products on}~C^{\infty}(V\times G)\llbracket\hbar\rrbracket\end{aligned}\Bigg\}$$\Bigg\{\text{star products on}~C^{\infty}(V\times G)\llbracket\hbar\rrbracket\Bigg\}\,.$}Alekseev-Calaque [Alekseev-Calaque] TheoremAXu [Xu-QuantumCNRS]Xu [Xu-Nonabelian]
@@
  • The left bottom-to-top-arrow is due to [Xu-Nonabelian]*Corollary 3.6.
  • The left top-to-bottom-arrow is pointed out in the proof of  [Alekseev-Calaque]*Proposition 1.3. The left vertical correspondence is one-to-one. In fact, this correspondence has appeared in the setting $\eta\subset\mathfrak{g}$ being abelian [Xu-R-matrices].
  • The right vertical correspondence is a bijection proved in [Xu-QuantumCNRS]. It is further shown in [Xu-QuantumCNRS] that star products can be equivalently described in terms of twistors of quantum groupoids.
@@
This note is structured to guide the reader through a comprehensive exploration of the topic. In Section 2, we revisit fundamental concepts related to Hopf algebroids, twistors, quantum groupoids, and dynamical twists. Section 3 is dedicated to stating and proving Theorem A. Finally, in Section 4, we apply Theorem A to the Enriquez-Etingof formal dynamical twists, which arise from nondegenerate polarized Lie algebras [Enriquez-Etingof2]. This application offers a novel perspective on equivariant star products on the associated Poisson homogeneous spaces, approached from the standpoint of quantum groupoids.
@@
[…]
@@ 2.1. Hopf algebroids and twistors
    $\Delta\colon{\mathcal{H}}\rightarrow{\mathcal{H}}\otimes_{R}{\mathcal{H}}$, called the coproduct, is an (R,R)(R\,,R)-bimodule morphism, satisfying $\Delta(1)=1\otimes 1$ and
@@
    $\epsilon\colon{\mathcal{H}}\rightarrow R$, called the counit, is an (R,R)(R\,,R)-bimodule map, satisfying $\epsilon(1)=1$ and
    $(\epsilon\otimes_{R}{\mathrm{id}})\circ\Delta=({\mathrm{id}}\otimes_{R}\epsilon)\circ\Delta={\mathrm{id}}.\,$
    Here we have used the identification $R\otimes_{R}{\mathcal{H}}\cong{\mathcal{H}}\otimes_{R}R\cong{\mathcal{H}}$.
@@
  • We follow [Lu] to add Condition (6) in our definition, which is however not required in [Xu-Quantum]. In fact, we need this condition to endow $R$ with a left $\mathcal{H}$}-module structure $\rhd\colon{\mathcal{H}}\times R\rightarrow R$ defined by
@@
[…]
@@ 2.2. Quantum groupoids and star products
[…]
@@
[…]
@@
The Hopf algebroid structure on $U(TM)$ over $C^{\infty}(M)$ extends naturally to a quantum groupoid structure on ${\mathcal{H}}:=U(TM)\llbracket\hbar\rrbracket$ over $R:=C^{\infty}(M)\llbracket\hbar\rrbracket$ . Consider a formal power series in $\hbar$
@@
for all $f,g\in C^{\infty}(M)\llbracket\hbar\rrbracket$, is a star product on $M$.
@@
[…]
$U({\mathfrak{l}}_{h})/\hbar U({\mathfrak{l}}_{h})\cong S({\mathfrak{l}})\cong\mathrm{Pol}({\mathfrak{l}}^{\ast}),$
@@
for all $f,g\in C^{\infty}({\mathfrak{l}}^{\ast})$, where $\{-,-\}_{\pi_{{\mathfrak{l}}^{\ast}}}$ is the Poisson bracket induced by the linear Poisson structure $\pi_{{\mathfrak{l}}^{\ast}}$ on $\mathfrak{l}$$^{\ast}$, and $B_{k}$’s are a family of bi-differential operators on $\mathfrak{l}$$^{\ast}$ (see [Dito, Kathotia, Gutt]).
@@
Let $\mathfrak{g}$} be a Lie algebra such that ${\mathfrak{l}}\subset{\mathfrak{g}}$ is a Lie subalgebra. Let $G$ be a Lie group integrating $\mathfrak{g}$} and $L\subset G$ be a Lie subgroup integrating ${\mathfrak{l}}\subset{\mathfrak{g}}$. For any $l\in{\mathfrak{l}}\subset{\mathfrak{g}}$, denote by $\overset{\rightarrow}{l}$ the left invariant vector field on $L$ or on $G$ generated by $l$.
@@
    for some $F=1\otimes 1\otimes 1+O(\hbar)\in C^{\infty}(V)\llbracket\hbar\rrbracket\otimes U({\mathfrak{g}})\otimes U({\mathfrak{g}})$.
@@
An $F$ subject to the equations (2.11) and (2.12) is called a smooth dynamical twist if it is also $\mathfrak{l}$}-invariant [Enriquez-Etingof2, Enriquez-Etingof1]. Thus, a right $L$-invariant compatible star product on $V\times G$ determines a smooth dynamical twist. For the converse, see Theorem 3.1 and the explanation on its relations with some existing works in the next section.